OpenStudy (yamyam70):

can someone explain this? 3 log 2x = 12 2x = 10^4 2x = 10 , 000 x= 5 000

OpenStudy (anonymous):

log 2x=4 2x=10^4

OpenStudy (yttrium):

3 log 2x = 12 You can simplify it by dividing the whole eq'n by 3. Hence, log 2x = 4 We all know that log (a) = log base10 (a) So therefore, it is the same as 10^4 = 2x 10000 = 2x x = 5000

OpenStudy (anonymous):

antilogging we get 10^4 as the base of log is 10

OpenStudy (yttrium):

did you get it @yamyam70 ?

OpenStudy (anonymous):

if the base were e,we would have got e^4

OpenStudy (anonymous):

have u unrstood the first step? log 2x=4 ???

OpenStudy (yamyam70):

one moment ,

OpenStudy (yamyam70):

log base (a) = 10 ?

OpenStudy (yamyam70):

log = 10

OpenStudy (anonymous):

base is 10

OpenStudy (yamyam70):

or log base 10 ?

OpenStudy (yttrium):

No.\[\log a = \log_{10} a \] i mean, if there is no base beside the log, it is automatically 10

OpenStudy (anonymous):

log(base 10)2x

OpenStudy (yamyam70):

so everytime I see " log " it automatically has base 10 ?

OpenStudy (yamyam70):

is my conclusion correct?

OpenStudy (yttrium):

yes :)

OpenStudy (yamyam70):

so , 10^4 = 2x why ?

OpenStudy (anonymous):

we r antilogging it,so log is removed and the base shifts to the other side

OpenStudy (anonymous):

if we have to find the value then we have to remove log first.

OpenStudy (anonymous):

to remove log we antilog it....and dis is the process.in dis sum we have to find the value of x.so its must be free frm all logs and coefficeints.

OpenStudy (yamyam70):

when you say , antilogging, we are simply converting it to an exponential equation right?

OpenStudy (unklerhaukus):

log 2x = 4 raise both sides to power 10 10^(log 2x) = 10^4 cancles 2x = 10^4

OpenStudy (anonymous):

yes u r right

OpenStudy (yttrium):

Do you remember the form \[\log_{a} b = c \] is same as \[a^c = b\] that's why it became 2x = 10^4

OpenStudy (yamyam70):

I get it now , thanks for the help everyone @Yttrium @UnkleRhaukus @madrockz :)

OpenStudy (anonymous):

u r welcome!!